# UniHENN: Designing More Versatile Homomorphic Encryption-based CNNs without im2col

### v1

@@ VI-B3 USPS Dataset
The USPS dataset [19] consists of 9,298 grayscale images of handwritten digits, each with dimensions $16\times 16$ pixels. The dataset also contains 10 classes, representing the digits from 0 to 9. The dataset consists of two parts: 7,291 data for training and 2,007 data for testing. We use these settings in our experiments.
@@ VI-B4 ECG Dataset
The ECG dataset from the MIT-BIH arrhythmia database [38] is used for the performance evaluation of our proposed patient-specific ECG approach. It has a total of 109,446 samples with a sampling frequency of 125Hz. The signals represent electrocardiogram (ECG) shapes of heartbeats in normal cases and cases affected by various arrhythmias and myocardial infarction. Each data point is labeled with one of 5 classes. In Abuadbba et al. [3], they pre-processed the ECG dataset to choose a total of 43 out of 48 records are selected and each signal was rescaled by the min-max normalization. And, choose the biorthogonal wavelet as decomposition and reconstruction wavelet function. We use 13,245 data for training and 13,245 data for testing which is in the https://github.com/SharifAbuadbba/split-learning-1D.
We use a preprocessed dataset generated by Abuadbba et al. [3], where each data point is one-dimensional and has 128 columns.
VI-B5 Hyperparameter Settings for Training
We trained all models with the learning rate of 0.001 and the batch size of 32. We trained the MNIST and USPS models for 15 epochs, the CIFAR-10 model for 25 epochs, and the ECG model for 30 epochs.
VI-C Library and Parameter Setup
We used SEAL-Python [2], a library that provides Python APIs for performing homomorphic operations like key generation, encryption, decryption, addition, multiplication, and rotation. The reason for using the SEAL-based library is that baseline models TenSEAL, and PyCrCNN in our experiments, are based on SEAL. So, we use it to balance the experiment. And, we use the pytorch and torchvision library for training. The detailed information of libraries and the experimental setup is described in the GitHub site : https://github.com/hm-choi/uni-henn. The specific parameters used in our experiments are detailed in Table II. All security parameters meet the 128-bit security level as described in [43].
TABLE II: Parameters used in UniHENN. $PK$ and $SK$ represent the size of the public key and the secret key. $GK$ represents the size of the Galois key, $RK$ represents the size of the set of Relinearization keys, and # mult means the total number of allowed multiplications.
@@ VI-C Library and Parameter Setup
The public key and secret key are also referred to as the encryption key and decryption key, respectively. The set consisting of the Galois key and the relinearization keys is also known as the evaluation key.
@@ VII Experimental Results
  • Comparison of Inference Time with State-of-the-Art Solutions. In this section, we compare the inference performance of UniHENN against two state-of-the-art HE-based deep learning inference frameworks: TenSEAL [7] and PyCrCNN [16]. We selected TenSEAL for comparison because it is a well-known open-source library that uses the im2col algorithm to implement CNNs. This algorithm allows TenSEAL to achieve highly optimized and efficient inference times for convolution operations. However, this efficiency comes at the cost of flexibility: TenSEAL’s architecture only supports models with a single convolutional layer. Therefore, TenSEAL is not a viable option for CNN architectures with multiple convolutional layers. PyCrCNN is an alternative to TenSEAL that does not employ the im2col algorithm, making it applicable to a broader range of CNN architectures. However, this flexibility comes with a significant drawback: PyCrCNN exhibits slower inference times than TenSEAL. Our experiments show that UniHENN strikes a balance between these extremes. It achieves inference times comparable to those of TenSEAL while also supporting CNN architectures with multiple convolutional layers. This makes UniHENN a more versatile solution for HE-based deep learning inference.
  • Adaptability of UniHENN for Various CNN Model Architectures. One of the compelling features of UniHENN is its ability to adapt to a variety of CNN architectures, including both complex and 1D CNN models. To demonstrate this adaptability, we conducted experiments with several CNN models, including LeNet-5, a seminal CNN model widely adopted for digit recognition tasks [51, 36]. LeNet-5 is more complex than many other models, with multiple convolutional layers and fully connected layers. Our successful implementation of LeNet-5 in UniHENN highlights the system’s ability to handle large and complex CNN models effectively. We also consider a 1D CNN architecture particularly useful for sequence-based tasks such as time-series analysis, natural language processing, and certain bioinformatics applications. This capability sets UniHENN apart from solutions like TenSEAL, which is constrained to supporting only single convolutional layer models and does not offer support for 1D CNNs. Through these experiments, we aim to demonstrate UniHENN’s comprehensive applicability and its ability to adapt to various CNN architectures, making it a versatile tool for secure and efficient deep learning inference across diverse application domains.
@@ VII-A Experiment 1: Comparison of Inference Time between UniHENN, TenSEAL, and PyCrCNN
In this experiment, we implement the model, denoted as $M_{1}$, using UniHENN, TenSEAL, and PyCrCNN. The $M_{1}$ model was introduced in the TenSEAL paper [7] and is well-suited for implementation using TenSEAL. Despite its simple architecture, which contains only one convolutional layer, the model achieves a high accuracy of 97.65%. Table III provides detailed specifications for this model architecture. The Conv2D parameters–$ch_{in}$, $ch_{out}$, size, and stride–represent the number of input channels, the number of output channels, kernel size, and stride for a two-dimensional convolutional layer, respectively. Similarly, $data_{in}$ and $data_{out}$ in the FC layer indicate the input and output dimensions. We use these notations to represent all other remaining models presented in this paper.
@@
Table IV presents the results. In terms of inference time for a single sample, the TenSEAL implementation exhibited the highest performance, taking 6.298 seconds. TenSEAL achieved this efficiency by optimizing convolution operations on the input data using the im2col algorithm. UniHENN took 16.248 seconds, making it 2.58 times slower than TenSEAL. PyCrCNN was the slowest, requiring 154.494 seconds. However, UniHENN has the advantage when handling multiple data points simultaneously. It outperforms TenSEAL by supporting batch operations for up to ten input samples, allowing concurrent inference calculations. TenSEAL and PyCrCNN do not offer parallel processing for multiple samples, resulting in total times that increase proportionally with the number of input samples. To process ten input samples together, UniHENN still takes 16.248 seconds, approximately 3.9 times and 94.6 times faster than TenSEAL’s 63.706 seconds and PyCrCNN’s 1537.228 seconds, respectively.
@@ VII-B Experiment 2: Comparison of Inference Time between UniHENN and PyCrCNN for LeNet-1
In this experiment, we utilize the LeNet-1 model, denoted as $M_{2}$. The model accuracy of $M_{2}$ is 98.62%, slightly higher than $M_{1}$. The detailed specifications of these model architectures can be found in Table V. Implementing the hyperbolic tangent function (tanh) in HE is computationally challenging; therefore, we have substituted the activation function from the tanh to the square function. We find that the modified LeNet-1 model with the square activation achieves an accuracy of 98.62% comparable to the 98.41% accuracy of the original LeNet-1 model when tested on a 10,000 sample MNIST dataset. This indicates that the modification in the activation function does not significantly impact the model’s accuracy.
@@
Total 794.064 (2.279) 30.090 (0.155)
In Table VI, UniHENN takes 30.090 seconds, making it approximately 26.4 times faster than PyCrCNN, which takes 794.064 seconds. In this experiment, most of the computational time for both UniHENN and PyCrCNN is consumed in the second convolutional layer. These findings highlight the importance of optimizing convolutional layer operations for time-efficient CNN inference in the context of HE.
@@ VII-C Experiment 3: Adaptibility of UniHENN for CNN Models with Approximate ReLU activation
Table VIII shows a layer-by-layer breakdown of the inference time for the $M_{3}$ model architecture. The Flatten and FC1 layers are the most time-consuming, taking an average of 9.603 seconds and 15.557 seconds, respectively. These two layers alone contribute significantly to the total inference time of 29.104 seconds. While UniHENN is significantly slower than non-HE models, it provides enhanced privacy and security, which may be crucial for specific applications or compliance requirements. The inference time of 29.104 seconds is still under 30 seconds, suggesting that UniHENN is practical for real-world applications, especially in contexts where data security is paramount. This inference time could be considered acceptable depending on the specific use case and the sensitivity of the data being processed. These results further validate the adaptability and efficiency of UniHENN in handling various CNN architectures with customized functionalities like approximate ReLU.
@@ VII-D Experiment 4: Adaptibility of UniHENN for LeNet-5
Table X shows a layer-by-layer breakdown of the inference time for the $M_{4}$ model architecture. The results show that the inference time for $M_{4}$ is 740.127 seconds. This is substantially slower than the previous experiments, which could be attributed to the model’s increased complexity. The slow performance emphasizes the need for further optimization, especially if UniHENN is to be broadly applied to more complex CNN architectures like LeNet-5 for real-world applications. Note that the convolutional layers are the most time-consuming, highlighting the critical need for optimizing these operations when implementing CNNs using HE. With such a long inference time, the immediate practicality of using this model for real-time or near-real-time applications is limited. However, this could be acceptable for services that require strong privacy controls and where data security is a higher priority than speed. For example, in healthcare or financial services, where data may be extremely sensitive, this level of privacy may justify the slower inference times.
@@ VII-E Experiment 5: Adaptability of UniHENN for CNN Models on Color Images
Table XII presents the experimental results, showing that the total inference time for UniHENN is approximately 20 minutes. While this may seem long, it is important to note that the convolutional layers are particularly time-consuming, requiring about 256.000 seconds for the first layer and 938.446 seconds for the second. This is consistent with the results of previous experiments. Despite the long inference time, we consider it tolerable given the inherent complexities of performing inference on color images, a capability not offered by alternative solutions.
@@
VII-F Experiment 6: Comparison of Inference Time of UniHENN on Grayscale Images
@@
Table XIV shows that UniHENN is approximately 6 times faster than PyCrCNN, with a total inference time of 12.307 seconds compared to PyCrCNN’s 71.139 seconds. Interestingly, the FC1 layer in PyCrCNN consumes a significant portion of the time (42.792 seconds), likely due to its large input size of 294. In contrast, although UniHENN also incurs a relatively high computational cost at the FC1 layer, it is substantially less than that of PyCrCNN. This suggests that UniHENN can efficiently handle FC layers with large input sizes.
@@
Total 71.139 (0.198) 12.307 (0.092)
VII-G Experiment 7: Comparison of Inference Time for a 1D CNN Model between UniHENN and PyCrCNN
@@
Table XVI shows that UniHENN is approximately 3.03 times faster than PyCrCNN, recording a total inference time of 5.119 seconds compared to PyCrCNN’s 15.513 seconds. This indicates that UniHENN is also more efficient even for relatively smaller models, such as 1D CNNs, where the computational time for convolutional layers is not as extensive. This efficiency positions UniHENN as an ideal choice for privacy-sensitive applications like disease diagnosis systems, aligning well with the principles of HE.
@@ VII-H Validation of inference results on encrypted data
We selected $M_{4}$ and $M_{5}$ as representative models for validation, as the errors were minimal in other models. We evaluated the $M_{4}$ model with 2,000 samples from the MNIST dataset and the $M_{5}$ model with 2,000 samples from the CIFAR-10 dataset. The results showed that all outputs produced by UniHENN were equivalent to those of the original models without any significant loss in accuracy. This indicates that UniHENN can perform secure and highly accurate inferences under the parameter configurations presented in Table II.
@@ VII-I Impact of HE Parameters
We conducted an additional analysis to investigate how the inference time and encrypted inference error vary depending on the parameters of the CKKS scheme, the HE algorithm we employ.
The CKKS scheme’s operation time and decrypted result accuracy are influenced by several parameters, specifically # slots, scale, and depth. Therefore, the inference time for CNN models is intricately linked to these parameters. A detailed analysis of these parameter settings is essential to optimize inference time while maintaining result accuracy.
The # slots parameter is determined by the degree $N$ of the polynomial ring, where $N$ is half the ring’s value. The scale parameter represents the precision of floating-point digits, while depth is influenced by both scale and the polynomial $P$.
@@ VII-I1 Influence of Depth on Inference Time
Figure 15: Average inference time (seconds) for $M_{1}$ with depth.

### v3

@@ VI-B3 USPS Dataset
The USPS dataset [21] contains 9,298 grayscale images of handwritten digits, each sized $16\times 16$ pixels, with 10 classes. We used 7,291 images for training and 2,007 images for testing.
@@ VI-B4 ECG Dataset
The ECG dataset from the MIT-BIH arrhythmia database [39] consists of 109,446 samples at a sampling frequency of 125 Hz. It includes signals representing ECG shapes of heartbeats in normal cases and various arrhythmias. Following the preprocessing steps by Abuadbba et al. [4], we used a subset of 26,490 samples, with 13,245 samples for training and 13,245 samples for testing, available at https://github.com/SharifAbuadbba/split-learning-1D.
VI-C Hyperparameter Settings for Training
We trained all models with a learning rate of 0.001 and a batch size of 32. The MNIST and USPS models were trained for 15 epochs, the CIFAR-10 model for 25 epochs, and the ECG model for 30 epochs, based on their convergence behavior.
VI-D Library and Parameter Setup
We used SEAL-Python [2] for homomorphic operations, ensuring consistency with our baseline models TenSEAL and PyCrCNN, which are also SEAL-based. PyTorch and torchvision were used for model training. Detailed information about libraries and the experimental setup is available at our GitHub repository: https://github.com/hm-choi/uni-henn. Table II provides the specific parameters used in our experiments. All security parameters comply with the 128-bit security level as described in [45].
TABLE II: Parameters used in UniHENN. $PK$: public key size (encryption key), $SK$: secret key size (decryption key), $GK$: Galois key size, $RK$: Relinearization keys size. $GK$ and $RK$ together form the evaluation key. # mult: total allowed multiplications.
@@ VII Experimental Results
  • Comparison of Inference Time with State-of-the-Art Solutions. In this section, we compare the inference performance of UniHENN against two state-of-the-art HE-based deep learning inference frameworks: TenSEAL [7] and PyCrCNN [17]. We selected TenSEAL for comparison because it is a well-known open-source library that uses the im2col algorithm to implement CNNs. This algorithm allows TenSEAL to achieve highly optimized and efficient inference times for convolution operations. However, this efficiency comes at the cost of flexibility: TenSEAL’s architecture only supports models with a single convolutional layer, making it unsuitable for CNN architectures with multiple convolutional layers. PyCrCNN, on the other hand, does not employ the im2col algorithm, making it applicable to a broader range of CNN architectures. However, this flexibility results in slower inference times compared to TenSEAL. Our experiments show that UniHENN successfully balances both flexibility and efficiency. It achieves inference times comparable to those of TenSEAL while supporting CNN architectures with multiple convolutional layers, making UniHENN a more versatile solution for HE-based deep learning inference.
  • Adaptability of UniHENN for Various CNN Model Architectures. One of the compelling features of UniHENN is its ability to adapt to a variety of CNN architectures, including both complex and 1D CNN models. To demonstrate this adaptability, we conducted experiments with several CNN models, including LeNet-5, a seminal CNN model widely adopted for digit recognition tasks [54, 37]. LeNet-5 is more complex than many other models, with multiple convolutional layers and fully connected layers. Our successful implementation of LeNet-5 in UniHENN highlights the system’s ability to handle large and complex CNN models effectively. We also considered a 1D CNN architecture particularly useful for sequence-based tasks such as time-series analysis, natural language processing, and certain bioinformatics applications. This capability sets UniHENN apart from solutions like TenSEAL, which is constrained to supporting only single convolutional layer models and does not offer support for 1D CNNs. Through these experiments, we aim to demonstrate UniHENN’s comprehensive applicability and its ability to adapt to various CNN architectures, making it a versatile tool for secure and efficient deep learning inference across diverse application domains.
@@ VII-A Experiment 1: Comparison of Inference Time between UniHENN, TenSEAL, and PyCrCNN
In this experiment, we implement the model, denoted as $M_{1}$, using UniHENN, TenSEAL, and PyCrCNN. The $M_{1}$ model was introduced in the TenSEAL paper [7] and is well-suited for implementation using TenSEAL. Despite its simple architecture, which contains only one convolutional layer, the model achieves a high accuracy of 97.65%. Table III provides detailed specifications for this model architecture. The Conv2D parameters – $CH_{in}$, $CH_{out}$, $K$, and $S$ – represent the number of input channels, the number of output channels, kernel size, and stride for a two-dimensional convolutional layer, respectively. Similarly, $DAT_{in}$ and $DAT_{out}$ in the FC layer indicate the input and output dimensions. We use these notations to represent all other remaining models presented in this paper.
@@
Table IV presents the results. In terms of inference time for a single sample, the TenSEAL implementation exhibited the highest performance, taking 6.298 seconds. TenSEAL achieved this efficiency by optimizing convolution operations on the input data using the im2col algorithm. UniHENN took 16.247 seconds, making it 2.6 times slower than TenSEAL. PyCrCNN was the slowest, requiring 154.494 seconds. However, UniHENN has the advantage when handling multiple data points simultaneously. It outperforms TenSEAL by supporting batch operations for up to ten input samples, allowing concurrent inference calculations. TenSEAL and PyCrCNN do not offer parallel processing for multiple samples, resulting in total times that increase proportionally with the number of input samples. To process ten input samples together, UniHENN still takes 16.247 seconds, approximately 3.9 times and 94.6 times faster than TenSEAL’s 63.706 seconds and PyCrCNN’s 1537.227 seconds, respectively.
@@ VII-B Experiment 2: Comparison of Inference Time between UniHENN and PyCrCNN for LeNet-1
In this experiment, we utilize the LeNet-1 model, denoted as $M_{2}$. The accuracy of $M_{2}$ is 98.62%, slightly higher than $M_{1}$. The detailed specifications of these model architectures can be found in Table V. Implementing the hyperbolic tangent function (tanh) in HE is computationally challenging; therefore, we have substituted the activation function from tanh to the square function. We find that the modified LeNet-1 model with the square activation achieves an accuracy of 98.62% comparable to the 98.41% accuracy of the original LeNet-1 model when tested on a 10,000 sample MNIST dataset. This indicates that the modification in the activation function does not significantly impact the model’s accuracy.
@@
Total 800.591 (2.279) 30.089 (0.155)
In Table VI, UniHENN takes 30.089 seconds, making it approximately 26.6 times faster than PyCrCNN, which takes 800.591 seconds. In this experiment, most of the computational time for both UniHENN and PyCrCNN is consumed in the second convolutional layer. These findings highlight the importance of optimizing convolutional layer operations for time-efficient CNN inference in the context of HE.
@@ VII-C Experiment 3: Adaptibility of UniHENN for CNN Models with Approximate ReLU activation
Table VIII shows a layer-by-layer breakdown of the inference time for the $M_{3}$ model architecture. The Flatten and FC1 layers are the most time-consuming, taking an average of 9.603 seconds and 15.557 seconds, respectively. These two layers alone contribute significantly to the total inference time of 29.105 seconds. While UniHENN is significantly slower than non-HE models, it provides enhanced privacy and security, which may be crucial for specific applications or compliance requirements. The inference time of 29.105 seconds is still under 30 seconds, suggesting that UniHENN is practical for real-world applications, especially in contexts where data security is paramount. This inference time could be considered acceptable depending on the specific use case and the sensitivity of the data being processed. These results further validate the adaptability and efficiency of UniHENN in handling various CNN architectures with customized functionalities like approximate ReLU.
@@ VII-D Experiment 4: Adaptibility of UniHENN for LeNet-5
Table X shows a layer-by-layer breakdown of the inference time for the $M_{4}$ model architecture. The results show that the inference time for $M_{4}$ is 740.128 seconds. This is substantially slower than the previous experiments, which could be attributed to the model’s increased complexity. The slow performance emphasizes the need for further optimization, especially if UniHENN is to be broadly applied to more complex CNN architectures like LeNet-5 for real-world applications. Note that the convolutional layers are the most time-consuming, highlighting the critical need for optimizing these operations when implementing CNNs using HE. With such a long inference time, the immediate practicality of using this model for real-time or near-real-time applications is limited. However, this could be acceptable for services that require strong privacy controls and where data security is a higher priority than speed. For example, in healthcare or financial services, where data may be extremely sensitive, this level of privacy may justify the slower inference times.
@@ VII-E Experiment 5: Adaptability of UniHENN for CNN Models on Color Images
Table XII presents the experimental results, showing that the total inference time for UniHENN is approximately 21 minutes. While this may seem long, it is important to note that the convolutional layers are particularly time-consuming, requiring about 256.000 seconds for the first layer and 938.446 seconds for the second. This is consistent with the results of previous experiments. Despite the long inference time, we consider it tolerable given the inherent complexities of performing inference on color images, a capability not offered by alternative solutions.
@@
VII-F Experiment 6: Evaluation of Inference Time of UniHENN on Grayscale Images
@@
Table XIV shows that UniHENN is approximately 6 times faster than PyCrCNN, with a total inference time of 12.309 seconds compared to PyCrCNN’s 71.139 seconds. Interestingly, the FC1 layer in PyCrCNN consumes a significant portion of the time (42.792 seconds), likely due to its large input size of 294. In contrast, although UniHENN also incurs a relatively high computational cost at the FC1 layer, it is substantially less than that of PyCrCNN. This suggests that UniHENN can efficiently handle FC layers with large input sizes.
@@
Total 71.139 (0.198) 12.309 (0.092)
VII-G Experiment 7: Evaluation of Inference Time of UniHENN for a 1D CNN Model
@@
Table XVI shows that UniHENN is approximately 3.0 times faster than PyCrCNN, recording a total inference time of 5.119 seconds compared to PyCrCNN’s 15.514 seconds. This indicates that UniHENN is also more efficient even for relatively smaller models, such as 1D CNNs, where the computational time for convolutional layers is not as extensive. This efficiency positions UniHENN as an ideal choice for privacy-sensitive applications like disease diagnosis systems, aligning well with the principles of HE.
@@ VII-H Validation of inference results on encrypted data
We selected $M_{4}$ and $M_{5}$ as representative models for validation, as the errors were very small in the case of other models. We evaluated the $M_{4}$ model with 2,000 samples from the MNIST dataset and the $M_{5}$ model with 2,000 samples from the CIFAR-10 dataset. The results showed that all outputs produced by UniHENN were equivalent to those of the original models without any significant loss in accuracy. This indicates that UniHENN can perform secure and highly accurate inferences under the parameter configurations presented in Table II.
@@ VII-I Impact of HE Parameters
We conducted an additional analysis to investigate how the inference time and encrypted inference error vary depending on the parameters of the CKKS scheme.
The CKKS scheme’s operation time and decrypted result accuracy are influenced by several parameters, specifically # slots, scale, and depth. Therefore, the inference time for CNN models is intricately linked to these parameters. A detailed analysis of these parameter settings is essential to optimize inference time while maintaining result accuracy. The # slots parameter is determined by the degree $N$ of the polynomial ring, where $N$ is half the ring’s value. The scale parameter represents the precision of floating-point digits, while depth is influenced by both scale and the polynomial $P$.
@@ VII-I1 Influence of Depth on Inference Time
Figure 14: Average inference time for $M_{1}$ with depth.
@@ VII-I2 Impact of Scale on Error
VII-J Comparison of UniHENN and DP-based CNN
This section compares our UniHENN framework with differential privacy (DP) for privacy-preserving CNNs, providing insights into the trade-offs between HE and DP-based approaches. We implemented DP-SGD [3] using Opacus [52].
We conducted experiments on the MNIST dataset with model $M_{4}$, comprising 60,000 training and 10,000 test images. The Adam optimizer was used with a learning rate of 0.001, training for 15 epochs with a batch size of 32. For the DP-SGD implementation, we set the maximum per gradient norm to 1.0 and used $\delta=0.00001$, satisfying the condition $\delta<1/60,000$ for the MNIST training set size. We varied $\epsilon$ from 5 to 0.1 during the 15-epoch training. While there is no definitive guideline for choosing $\epsilon$, NIST’s post [41] suggests that $\epsilon$ between 0 and 5 provides strong privacy protection. Thus, our selected $\epsilon$ values with $\delta=0.00001$ would represent suitable candidates for DP-SGD-based training.
Tables XVII and XVIII present the accuracy and memory consumption results, respectively.
TABLE XVII: Accuracy comparison between UniHENN (U) and DP-based CNN (D) for various privacy budgets ($\epsilon$) (30 runs, standard deviations in parentheses).
Framework U D ($\epsilon$=5) D ($\epsilon$=1) D ($\epsilon$=0.5) D ($\epsilon$=0.1)
Accuracy (%) 98.91 84.19 (2.59) 84.57 (2.29) 83.73 (1.98) 80.63 (1.54)
TABLE XVIII: Memory consumption comparison during inference (30 runs, standard deviations in parentheses).
Framework U D
Memory (MiB) 3677.64 (1.05) 3.91 (0.10)
Our results reveal significant trade-offs between HE-based and DP-based privacy-preserving techniques in CNNs. UniHENN achieves superior accuracy (98.91%) compared to the DP-based CNN, which ranges from 84.57% to 80.63% as $\epsilon$ decreases. This demonstrates UniHENN’s ability to maintain model performance across varying privacy parameters. However, this high accuracy comes at a substantial computational cost: UniHENN consumes about 941 times more memory (3677.64 MiB) than the DP-based approach (3.91 MiB) due to the intensive nature of HE operations.
These findings highlight the critical resource trade-offs in privacy-preserving machine learning, emphasizing the need to balance accuracy and computational efficiency when selecting privacy-preserving techniques for specific applications. The choice between HE and DP approaches will depend on the particular use case, with UniHENN offering superior accuracy at the cost of higher computational resources, while DP-based methods provide a more memory-efficient solution with a trade-off in accuracy.